# Nighttime Person Re-Identification via Collaborative Enhancement Network with Multi-domain Learning

### v1

@@ 3.3 Relighting Branch
As shown in Figure 2 (a), the relighting branch comprises a shared encoder and a relighting subnet. Considering that image relighting is an image reconstruction task rather than a classification task, we exclude the $X_{cls}$ component from $F_{share}$ and denote it as $F_{share}^{*}$. Subsequently, we forward it to the relighting subnet to perform image relighting.
@@
Relighting Subnet. The relighting branch incorporates 6 Transformer decoders to decode the shared features $F_{share}^{*}$ and utilize them for subsequent image reconstruction. These decoders consist of multi-head self-attention modules and feedforward networks. Within each decoder layer, the attention calculation is performed using the global representation $F_{share}^{*}$ and the task-specific learnable embedding $T_{e}$ as inputs, as illustrated in Figure 2 (b). The output features of the decoder are further processed through a sequence of convolutional layers, combined with the maximum channel of the input image, to generate a reflectance map (relighting result) of the input image and a corresponding lighting map. Specifically, the output features of the decoder and the maximum channel of the input image undergo separate $3\times 3$ convolutional operations for refinement. Subsequently, they are concatenated along the channel dimension and passed through two additional convolutional layers with activation functions to enable comprehensive fusion, resulting in a 4-channel image. Finally, the Sigmoid function is applied to constrain these results. The first three channels represent reflection maps, while the last channel corresponds to the lighting map.
3.4 Multilevel Feature Interactions
@@
To improve ReID performance in low-light night scenes, we abandon the conventional approach [18, 36, 4] of extracting ReID features from relighting images. Instead, we focus on enhancing ReID feature representation from collaborative modeling between two tasks. Specifically, we adopt two strategies to achieve the enhancement from the relighting branch to the ReID branch. Since low-level networks typically capture common image features, we utilize parameter-sharing scheme between the ReID and relighting branches to improve low-level feature representation. Considering different tasks within high-level networks often model task-specific knowledge, which poses challenges for direct interaction. To this end, we propose a feature distillation strategy to enhance high-level feature in the ReID branch.
@@
Shared Encoder. The shared encoder is designed to extract crucial information from input nighttime images, serving both the joint learning ReID task and the relighting task. Unlike traditional convolutional neural networks with limited receptive fields, we equip this extractor with ViT encoders [6], which allows for better capturing of global information within nighttime images. Moreover, the extractor excludes downsampling operations, thus preserving the spatial structure and detailed information in shallower layers. This preservation is particularly advantageous for the image relighting process. Therefore, benefiting from the learning of the relighting task, the shared encoder will share the ability to extract lighting-enhanced features in both tasks.
@@
Feature Distillation. High-level networks tend to model task-specific knowledge in various tasks. Therefore, it is a challenging problem to leverage the high-level features of the relighting subnet to enhance the high-level features of the ReID subnet. To address this issue, we adopt a non-intrusive interactive strategy known as feature distillation to enhance ReID features. By analyzing the correlation between the two tasks, we introduce a content-independent brightness loss to assist the ReID subnet. However, this overall brightness enhancement might amplify the brightness differences in person captured from various viewpoints. Consequently, we also incorporate a feature contrastive loss to ensure consistent feature representation for person from the same class, even under different lighting conditions. Finally, based on the above losses, we propose a novel lighting distillation loss to achieve high-level feature interaction between relighting and ReID tasks and improve the performance of the ReID task. Specifically, the lighting distillation loss can be formulated as:
@@
$L_{LD}=\underbrace{||\varphi(f_{s}),\varphi(f_{t})||_{2}}_{\textit{brightness loss}}-\underbrace{log(\frac{exp(sim(f_{s},f_{t}))}{\frac{1}{N}{\textstyle\sum_{i=1}^{N}exp(sim(f_{s},f^{i}_{t})})}}_{\textit{contrastive loss}}$
@@
where $\varphi()$ represents the feature channel mean function, and $||\cdot||_{2}$ denotes a mean square function. $N$ is the number of samples in the dataset, $\text{sim}(u,v)$ denotes a similarity measure between $u$ and $v$, and $i$ and $j$ are indices of data points, indicating whether $x_{i}$ and $x_{j}$ are samples from the same class or different classes.
3.5 Multi Domain Learning
To overcome the challenge of the small size of existing nighttime ReID dataset, we synthesize a large-scale nighttime person ReID dataset by applying multi-dimensional random degradation to existing daytime person ReID datasets (Market1501 [42] and MSMT17 [28]). Specifically, we apply lighting, contrast, color, and hue as four degradation factors to daytime images, generating diverse nighttime images that correspond to real nighttime scenarios, as depicted in Figure 3.
Figure 3: Comparison of real and synthetic domain data samples
However, there are large inter-domain differences between synthetic data and real nighttime data. To tackle this challenge, this study introduces a novel multi-domain learning scheme, illustrated in Figure 4. Within this scheme, the ReID branch performs shared learning in the two domains, effectively enhancing its learning of domain-invariant knowledge. The learning process of the entire network alternates between the two domains, employing supervised relighting learning in the synthetic domain and unsupervised relighting learning in the real domain. In the following sections, we will provide a detailed explanation of the learning approaches adopted in both domains.
Figure 4: The overview of the proposed multi-domain learning.
@@ 3.5 Multi Domain Learning
Synthetic Domain Learning: In this stage, we establish a supervised low-light image relighting scheme utilizing synthetic low-light data and their corresponding well-lit counterparts. Given access to the corresponding ground truth, we employ the Euclidean distance as the loss function for this scheme, denoted as $\mathcal{L}_{IE_{s}}$, which can be expressed as:
$\mathcal{L}_{IE_{s}}=\frac{1}{N}\sum_{i=1}^{N}\|I_{L_{i}}-I_{H_{i}}\|^{2}_{2}$ (4)
@@
where $N$ represents the total number of training samples, $I_{L_{i}}$ and $I_{H_{i}}$ denote the synthetic low-light image and corresponding well-lit image, respectively. $||\cdot||_{2}$ indicates the Euclidean distance between two images. This loss function quantifies the discrepancy between the synthetic low-light image and its corresponding well-lit image, with the objective of minimizing this difference during training.
@@
$\mathcal{L}_{Syn}=\mathcal{L}_{ID}+\mathcal{L}_{Tri}+\lambda_{1}\mathcal{L}_{IE_{s}}+\lambda_{2}\mathcal{L}_{lig}$ (5)
@@
Real Domain Learning. Given the challenges of obtaining highly corresponding low-light person images at night and well-lit person images during the day in real-world scenes, we adopt an unsupervised relighting learning strategy to enhance low-light images. Specifically, we employ the RetineX model to decompose an image as follows:
$\vspace{-3pt}\mathcal{M}=\mathcal{R}\times\mathcal{I}.$ (6)
@@
Relighting in unsupervised scenarios presents a highly ill-posed problem. To address this issue, we draw inspiration from [37] and incorporate essential prior information along with four loss functions: (1) reconstruction loss $\mathcal{L}_{rec}$, (2) reflection loss $\mathcal{L}_{ref}$, (3) Color loss $\mathcal{L}_{col}$ (4) structure-aware smoothing loss $\mathcal{L}_{sa}$.
First, we utilize the reconstruction loss ($\mathcal{L}_{rec}$) to perserve the invariance of image content and structural information. In this study, we employ a combination of mean absolute error loss and structural similarity loss for jointly constrain:
$\mathcal{L}_{rec}=1-SSIM(\mathcal{R}\cdot\mathcal{I},\mathcal{M})+\|\mathcal{R}\cdot\mathcal{I}-\mathcal{M}\|_{1}$ (7)
[…]
@@
$\vspace{-6pt}|\Delta\mathcal{R}|=|\nabla_{x}\mathcal{R}|1+|\nabla{y}\mathcal{R}|_{1}$ (8)
$\Delta\mathcal{R}$ represents the difference between neighboring pixels in the reflection map, while $\nabla_{x}$ and $\nabla_{y}$ denote the gradients in the horizontal and vertical directions, respectively. Thus, the overall reflection loss can be defined as follows:
$\mathcal{L}_{ref}=\left\|\max_{c\in r,g,b}R^{c}-\mathcal{H}\left(\max_{c\in r,g,b}M^{c}\right)\right\|_{1}+||\Delta\mathcal{R}||_{1}$ (9)
@@
Third, a color loss is introduced to control the color shift problem during the enhancement process using the Gray-world color constancy hypothesis [3], which can be described as follows:
$\mathcal{L}_{col}=\sum_{\forall(i,j)\in\varepsilon}\left(R^{i}-R^{j}\right)^{2},\varepsilon=\{(r,g),(r,b),(g,b)\}.$ (10)
@@
$\vspace{-6pt}\mathcal{L}_{sa}=\|\Delta\mathcal{I}\cdot\exp(-\Delta\mathcal{R})\|_{1}.$ (11)
@@
$\vspace{-6pt}\mathcal{L}_{IE_{r}}=\lambda_{a}\mathcal{L}_{rec}+\lambda_{b}\mathcal{L}_{ref}+\lambda_{c}\mathcal{L}_{col}+\lambda_{d}\mathcal{L}_{sa}.$ (12)
In this study, we set $\lambda_{a}$, $\lambda_{b}$, $\lambda_{c}$, and $\lambda_{d}$ to 1, 0.1, 0.2, and 0.1, respectively. Finally, the overall learning loss function for real domain data is defined as:
$\vspace{-6pt}\mathcal{L}_{Real}=\mathcal{L}_{ID}+\mathcal{L}_{Tri}+\lambda_{1}\mathcal{L}_{IE_{r}}+\lambda_{2}\mathcal{L}_{lig}.$ (13)
Here, $\lambda_{1}$ and $\lambda_{2}$ are consistent with the settings of $\lambda_{1}$ and $\lambda_{2}$ setting in synthetic domain learning.
4 Experiment
In this section, we present an overview of our experimental setup and proceed to conduct a comprehensive evaluation and comparison of the proposed method on three datasets: two real nighttime ReID datasets and one synthetic ReID dataset. We follow the common evaluation protocol in the field of ReID. We employ two standard metrics: the Cumulative Matching Characteristic (CMC) curve and the mean Average Precision (mAP).
4.1 Experiment Settings
@@
Implementation Details. In this paper, CENet utilizes the TransReID [9] network as the backbone in the ReID branch. The backbone is equipped with an ImageNet pre-trained model for feature extraction. Corresponding classification head networks are designed for both the real domain dataset and the synthetic domain dataset, tailored to their respective number of categories. In the relighting branch, the shared encoder follows the aforementioned parameters. However, the relighting subnet uses randomly initialized parameters, which allows the framework to learn specific features for low-light image relighting. The entire framework is trained end-to-end on the real domain dataset Night600 and the synthetic domain dataset Syn_Dark. Specifically, all images are resized to a size of $256\times 128$, and random horizontal flipping, cropping, and random erasing techniques are applied during training to mitigate overfitting issues. Each batch size is set to 64, and the initial learning rate is set to 0.008. We employ stochastic gradient descent (SGD) optimizer with a momentum of 0.9 and weight decay of 1e-4. In testing phase, CENet can only perform the ReID branch, which gets rid of the extra computational cost of existing methods that rely on relighting results during inference.
[…]
Table 1: Statistics on real and synthetic domain datasets
@@
Table 2: Comparison results of CENet with state-of-the-art ReID methods on real and synthetic nighttime ReID datasets
@@
4.2 Comparison with State of the Arts

### v2

@@ III-C Relighting Branch
As shown in Figure 2, the relighting branch comprises a shared encoder and a relighting subnet. Considering that image relighting is an image reconstruction task rather than a classification task, we exclude the $X_{cls}$ component from $F_{share}$ and denote it as $F_{share}^{*}$. Subsequently, we forward it to the relighting subnet to perform image relighting.
Figure 3: The illustration details the transformer decoder, where $\bigoplus$ represents feature summation.
Most of the existing deep relighting studies [6, 10] are based on CNN network, which lacks the modeling ability of global lighting. To this end, we design a relighting branch based on Transformer with powerful global relationship modeling capabilities. The entire relighting branch contains a set of Transformer encoders fully shared with the ReID network and an independent relighting subnet. The relighting subnet incorporates six Transformer decoders to decode the shared features $F_{share}^{*}$ and utilize them for subsequent image reconstruction. These decoders consist of multi-head self-attention modules and feedforward networks. Unlike the original Transformer decoder, which performs decoding iteratively, we decode the features of each position in parallel. Thus, within each decoder layer, the attention calculation uses the global representation $F_{share}^{*}$ and the task-specific learnable embedding $T_{e}$, which serve as position embeddings. As shown in Figure 3, the computations are defined as follows:
$Q=W_{q1}(F_{share}^{*}+T_{e}),$ (4)
$K=W_{k1}(F_{share}^{*}+T_{e}),$
$V=W_{v1}(F_{share}^{*}),$
$D_{out}^{*}=Norm(SA(Q,K,V)+F_{share}^{*}),$
$Q^{*}=W_{q2}(D_{out}^{*}+T_{e}),$
$K^{*}=W_{k2}(F_{share}^{*}),$
$V^{*}=W_{v2}(F_{share}^{*}),$
$D^{**}_{out}=Norm(SA(Q^{*},K^{*},V^{*})+D_{out}^{*}),$
$D_{out}=Norm(FFN(D^{**}_{out})+D^{**}_{out}),$
where $W_{q1},W_{k1},W_{v1},W_{q2},W_{k2},W_{v2}$ represent the projection layers. $Norm(.),SA(.)$ and $FFN(.)$ refer to layer normalization, self-attention layer, and feed-forward network in that order. $D_{out}$ denotes the output feature of the decoder layer.
@@
The output features of the decoder are further processed through a sequence of convolutional layers, combined with the maximum channel of the input image, to generate a reflectance map (relighting result) of the input image and a corresponding lighting map. Specifically, the output features of the decoder and the maximum channel of the input image undergo separate $3\times 3$ convolutional operations for refinement. Subsequently, they are concatenated along the channel dimension and passed through two additional convolutional layers with activation functions to enable comprehensive fusion, resulting in a 4-channel image. Finally, the Sigmoid function is applied to constrain these results. The first three channels represent reflection maps, while the last channel corresponds to the lighting map. It is worth noting that we tailor different learning strategies for the relighting branch in both real and synthetic domain data. The details will be elaborated in Sec III-E.
III-D Multilevel Feature Interaction
@@
To improve ReID performance in low-light night scenes, we abandon the conventional approach [7, 8, 50] that extracts ReID features from relighting images. Instead, we focus on enhancing ReID feature representation from collaborative modeling between two tasks. Specifically, we adopt two strategies to achieve the enhancement from the relighting branch to the ReID branch. Since low-level networks typically capture common image features, we utilize parameter-sharing scheme between the ReID and relighting branches to improve low-level feature representation. Considering different tasks within high-level networks often model task-specific knowledge, which poses challenges for feature interaction. To handle this problem, we propose a feature distillation strategy to enhance high-level feature in the ReID branch.
@@
Low-level Shared Encoder. The shared encoder is designed to extract crucial information from input nighttime images, serving both the joint learning ReID task and the relighting task. Unlike traditional convolutional neural networks with limited receptive fields, we equip with ViT encoders [49], which allows for better capturing of global information in nighttime images. Moreover, we exclude downsampling operations, thus preserving the spatial structure and detailed information in the shallower layers. This preservation is particularly advantageous for the image relighting process. Therefore, benefiting from the learning of the relighting task, the shared encoder will share the ability to extract lighting-enhanced features in both tasks.
High-level Feature Distillation. High-level layers tend to model task-specific knowledge in various tasks. Therefore, it is a challenging problem to leverage the high-level features of the relighting subnet to enhance the high-level features of the ReID subnet. To address this issue, we adopt a non-intrusive interactive strategy known as feature distillation to enhance ReID features. By analyzing the correlation between the two tasks, we can conclude that the ReID task aims to improve the discriminative properties of foreground targets in an image, while the relighting task seeks to improve the global lighting level of an image. Therefore, there is some correlation between the two tasks. To achieve collaborative enhancement between the two tasks, we do not consider introducing an additional front-background separation technique for accurate distillation, instead utilizing a simple brightness loss function that is content-independent to boost the global lighting level of the ReID features.
@@
However, this overall brightness enhancement might amplify the brightness differences in person captured from various viewpoints. Consequently, we also incorporate a feature contrastive loss to ensure consistent feature representation for person from the same class, even under different lighting conditions. Finally, based on the above losses, we propose a novel feature distillation loss to achieve high-level feature interactions between relighting and Re-ID tasks and improve the performance of the ReID task. Specifically, the feature distillation loss consists of two items, the first of which aims to make the brightness features in ReID network approximate to the light-enhanced features. However, we observe that the performance improvement is limited when only using brightness loss distillation, as shown in Table LABEL:tab:tcte_ablation2. We argue that although the enhancement of feature brightness can improve the feature representation ability, the risk of possible over-enhancement tends to decrease the differences between different individuals. Therefore, we introduce a contrast learning loss [51] to pull together the features of persons in the same category with different brightnesses and push away the features of persons in different categories with similar brightness. Overall, the feature distillation loss can be formulated as:
@@
$\mathcal{L}_{FD}=\underbrace{||\varphi(f_{s}),\varphi(f_{t})||_{2}}_{\textit{brightness loss}}-\underbrace{log(\frac{exp(sim(f_{s},f_{t}))}{\frac{1}{N}{\textstyle\sum_{i=1}^{N}exp(sim(f_{s},f^{i}_{t})})})}_{\textit{contrastive loss}},$ (5)
@@
where $f_{s}$ and $f_{t}$ denote the high-level features of the two tasks, respectively. $\varphi()$ represents the feature channel mean function, and $||\cdot||_{2}$ denotes a mean square function. $N$ is the number of samples in the dataset, $\text{sim}(u,v)$ denotes a similarity measure between $u$ and $v$, and $i$ and $j$ are indices of data points, indicating whether $x_{i}$ and $x_{j}$ are samples from the same classes or different classes.
III-E Multi-Domain Learning
Figure 4: Comparison of real and synthetic domain data samples.
Constrained by the annotation difficulty of nighttime images, existing nighttime ReID datasets are small scale, which is hard to train big networks such as Transformer. To overcome these challenges, we synthesize a large-scale nighttime person ReID dataset by applying multi-dimensional random degradation to existing daytime person ReID datasets (Market1501 [2] and MSMT17 [3]). As depicted in Figure 4, we apply lighting, contrast, color, and hue as four degradation factors to daytime images, generating diverse nighttime images that correspond to real nighttime scenarios. This approach offers two distinct advantages. First, the synthesized nighttime images maintain a one-to-one correspondence with their daytime counterparts, facilitating the establishment of supervised relighting learning strategies. Second, the introduction of synthetic data significantly enhances the diversity within the training set, thereby playing a crucial role in improving the model’s generalization capabilities. As shown in Figure 5, the ReID branch performs shared learning in the two domains, effectively enhancing its learning of domain-invariant knowledge. The learning process of the entire network alternates between the two domains, employing supervised relighting learning in the synthetic domain and unsupervised relighting learning in the real domain. In the following sections, we will provide a detailed explanation of the learning approaches adopted in both domains.
Figure 5: Overview of the proposed multi-domain learning approach.
@@ III-E Multi-Domain Learning
Synthetic Domain Learning. In this stage, we establish a supervised low-light image relighting scheme utilizing synthetic low-light data and their corresponding well-lit counterparts. Given access to the corresponding ground truth, we employ the Euclidean distance as the loss function for this scheme, denoted as $\mathcal{L}_{IE_{s}}$, which can be expressed as:
$\mathcal{L}_{IE_{s}}=\frac{1}{N}\sum_{i=1}^{N}\|I_{L_{i}}-I_{H_{i}}\|^{2}_{2},$ (6)
@@
where $N$ represents the total number of training samples, $I_{L_{i}}$ and $I_{H_{i}}$ denote the relighting result of the synthetic low-light image and corresponding well-lit image, respectively. $||\cdot||_{2}$ indicates a mean square function. This loss function quantifies the difference between the relighting result of the synthetic low-light image and its corresponding well-lit image, with the objective of minimizing this difference during training.
@@
$\mathcal{L}_{Syn}=\mathcal{L}_{ID}+\mathcal{L}_{Tri}+\lambda_{1}\mathcal{L}_{IE_{s}}+\lambda_{2}\mathcal{L}_{FD},$ (7)
@@
Real Domain Learning. Given the challenges of obtaining highly corresponding low-light person images at night and well-lit person images during the day in real-world scenes, we design an unsupervised relighting scheme to train the relighting network in real domain data. Specifically, we employ the RetineX model to decompose an image as follows:
$\mathcal{M}=\mathcal{R}\cdot\mathcal{I}.$ (8)
@@
Relighting in unsupervised scenarios presents a highly ill-posed problem. To address this issue, we draw inspiration from [10] and incorporate essential prior information along with four loss functions, including (1) reconstruction loss $\mathcal{L}_{rec}$, (2) reflection loss $\mathcal{L}_{ref}$, (3) color loss $\mathcal{L}_{col}$ (4) structure-aware smoothing loss $\mathcal{L}_{sa}$.
First, we utilize the reconstruction loss ($\mathcal{L}_{rec}$) to preserve the invariance of image content and structural information. In this study, we employ a combination of mean absolute error loss and structural similarity loss for jointly constrain:
$\mathcal{L}_{rec}=1-SSIM(\mathcal{R}\cdot\mathcal{I},\mathcal{M})+\|\mathcal{R}\cdot\mathcal{I}-\mathcal{M}\|_{1},$ (9)
where ${SSIM(u,v)}$ and $\|\cdot\|_{1}$ represent the structural similarity index and $l_{1}$ norm, respectively.
@@
$\Delta\mathcal{R}=||\nabla_{x}\mathcal{R}||_{1}+||\nabla{y}\mathcal{R}||_{1},$ (10)
where $\Delta\mathcal{R}$ represents the difference between neighboring pixels in the reflection map, while $\nabla_{x}$ and $\nabla_{y}$ denote the gradients in the horizontal and vertical directions, respectively. Thus, the overall reflection loss can be defined as follows:
$\mathcal{L}_{ref}=\left\|\max_{c\in r,g,b}R^{c}-\mathcal{H}\left(\max_{c\in r,g,b}M^{c}\right)\right\|_{1}+\Delta\mathcal{R},$ (11)
@@
Third, a color loss is introduced to mitigate color shifts during the enhancement process, leveraging the Gray-world color constancy hypothesis [52], which can be described as follows:
$\mathcal{L}_{col}=\sum_{\forall(i,j)\in\varepsilon}\left(R^{i}-R^{j}\right)^{2},\varepsilon=\{(r,g),(r,b),(g,b)\},$ (12)
@@
$\mathcal{L}_{sa}=\|\Delta\mathcal{I}\cdot\exp(-\Delta\mathcal{R})\|_{1}.$ (13)
@@
$\mathcal{L}_{IE_{r}}=\lambda_{a}\mathcal{L}_{rec}+\lambda_{b}\mathcal{L}_{ref}+\lambda_{c}\mathcal{L}_{col}+\lambda_{d}\mathcal{L}_{sa}.$ (14)
In this study, we follow [10] to set $\lambda_{a}$, $\lambda_{b}$, $\lambda_{c}$, and $\lambda_{d}$ to 1, 0.1, 0.2, and 0.1, respectively.
Finally, we employ the unsupervised relighting loss defined in Eq (14) and ReID loss to jointly optimize the whole network. Therefore, the overall loss function for real domain data is defined as:
$\mathcal{L}_{Real}=\mathcal{L}_{ID}+\mathcal{L}_{Tri}+\lambda_{1}\mathcal{L}_{IE_{r}}+\lambda_{2}\mathcal{L}_{FD}.$ (15)
Here, $\lambda_{1}$ and $\lambda_{2}$ are consistent with the settings of $\lambda_{1}$ and $\lambda_{2}$ setting in Eq (7).
IV Experiment
In this section, we present the experimental setup and the comprehensive evaluation and comparison of the proposed method on three datasets, including two real nighttime ReID datasets and one synthetic ReID dataset. In addition, we perform an ablation study to justify the effectiveness of each component in the proposed method.
IV-A Experiment Settings
@@
Implementation Details. In this work, we utilize the TransReID [4] network as the backbone of ReID branch. The backbone is equipped with an ImageNet pre-trained model for feature extraction. Corresponding classification heads are designed for both the real domain dataset and the synthetic domain dataset, tailored to their respective number of categories. In relighting branch, the shared encoder follows the aforementioned parameters. However, the relighting subnet uses randomly initialized parameters, which allow the network to learn specific features for image relighting. The entire framework is trained end-to-end on the real domain dataset Night600 and the synthetic domain dataset Syn_Dark. Specifically, all images are resized to a size of $256\times 128$, and random horizontal flipping, cropping, and random erasing techniques are applied during training to avoid overfitting. Each batch size is set to 64, and the initial learning rate is set to 0.008. We employ the stochastic gradient descent (SGD) optimizer with a momentum of 0.9 and weight decay of 1e-4. During the testing phase, CENet executes only the ReID branch, thereby eliminating the extra computational cost of image relighting in the inference stage.
Datasets. We utilize the Night600 dataset [7] as our real domain data. This dataset comprises a total of 28,813 nighttime images featuring 600 different individuals. The images are captured using 8 cameras and divided into a training set and a testing set. Specifically, the training set contains data from 300 individuals, while the testing set includes the remaining 300 individuals from the dataset.
We create a synthetic ReID dataset called Syn_Dark by adjusting the brightness, contrast, hue, and color parameter ranges of images from Market1501 [2] and MSMT17 [3] datasets. To simulate lighting conditions and color variations in nighttime images, we leverage professional image editing software, Photoshop V23.3.2, to create synthetic data. We adjust the input images in four dimensions, including brightness, contrast, hue, and color. To reduce brightness and contrast, we set the brightness parameter range between (10, 38) and the contrast parameter range between (7, 30). In addition, we set the image hue and color parameter ranges to (7, 30) and (20, 35), respectively. By randomly selecting a specific value within a predefined range, we apply the corresponding degradation proportion to the image. For example, if the randomly chosen parameter for brightness is 10, the image’s brightness will be reduced to 10% of its original level. The primary challenges of real nighttime image are low light, low contrast, and color shifts that occur under different low-light conditions. Therefore, we perform the degradation in the above four dimensions to maximally simulate the real data. By leveraging the random combinations of degradation parameters and the availability of large-scale daytime data, we are able to obtain a broader range of challenging nighttime images. Representative samples from both datasets are shown in Figure 4, highlighting differences in sharpness and noise between the real and synthetic domains.
TABLE I: Statistics of real and synthetic domain datasets.
@@
TABLE II: Comparison results of CENet with state-of-the-art ReID methods on real and synthetic nighttime ReID datasets.
@@
[…]
IV-B Comparison with State-of-the-art Methods